You are a population geneticist studying Gene B in a population that is in Hardy-Weinberg equilibrium. You sample a population to determine the genotypes at Gene B. You find that the freuency of homozygous recessive genotypes (bb) is 0.10. What would be the frequency of the dominant allele (B) in this population?

A. 0.84

B. 0.4

C. 0.36

D. 0.6

E. 0.48
Answer #1

The frequency of the dominant allele (B) in this population is............................0.68 or 68%

  • homozygous recessive genotype................bb = 0.10 or 10%
  • The genotype frequency of bb is 10% , that is q2 = 10%
  • q2 = 10% or 0.1
  • then q = 0.31
  • then the allelic frequency of recessive homozyogus "a" is .........31%
  • if q = 0.31
  • p +q =1, then p =1- q
  • p = 1-0.31 = 0.69
  • p = 69%
  • if p = 69%, q = 31%
  • then p2+2pq+q2 = 1
  • (0.69)2 + 2 (0.69)(0.31) + (0.31)2 = 1
  • 0.4761+0.4278+ 0.0961 =1
  • The genotypic frequency of homozygous dominant allele = p2 =47.6%
  • then p = p2+1/2 (2pq) = 0.476+1/2(0.4278) = 0.6899
  • p or dominant allele "B" phenotypic frequency = 0.6899
  • then q equals to 1-p = 1-0.6899 = 0.31
  • 68 % of the population show dominant phenotype.
  • 31 % of the population show recessive phenotype.